OpenStudy (kitten_is_back):

Column B 1. Four more than 6 times a number is 34. 2. The difference of 4 times a number and 6 is 34. 3. Six more than 4 times a number is 34. 4. The sum of 4 times a number and 6 is 34. A. 4n + 6 = 34 B. 4n – 6 = 34 C. 6n + 4 = 34

OpenStudy (anonymous):

1. is C. 6n + 4 = 34 since 4 plus 6 times a number is really 6 times a number plus 4

OpenStudy (kitten_is_back):

ohh

OpenStudy (anonymous):

3. B Six more than 4 times a number is 34. since 4n+6=34 when n represents "a number"

OpenStudy (kitten_is_back):

okay

OpenStudy (kitten_is_back):

2 a ?

OpenStudy (anonymous):

3 is A

OpenStudy (anonymous):

sorry 'bout "3. B Six more than 4 times a number is 34. since 4n+6=34 when n represents "a number"" 3 is actually A